OpenStudy (anonymous):

Show the expanded form and simplification of x^6/x^2. Explain in your own words how you can simplify x^6/x^2 without having to write the expanded form.

OpenStudy (oaktree):

Do you mean \[\frac{ x^6 }{ x^2 }\]Or \[(\frac{ x^6 }{ x })^2\]

OpenStudy (anonymous):

the first one

OpenStudy (oaktree):

Okay. So it's just \[x^{6-2} = x^4 = x \times x \times x \times x\]And we can write it out as just x^4 because that's what exponentiation does for mathematicians.